A bar of metal contains 10% silver. Another bar contains 15% silver. How many kilograms of each must be taken to make a 10 kilogram bar containing 12% silver.
Answer by RAY100(1637)

Let x = amt at 10%,,,and y= amt at 15%
.
(1) x + y = 10,,,,,,,or y=10-x
.
(2) .10*x +.15*y = .12*10,,,,,,,,,silver balance
.
subst from (1)
.
.1x +.15(10-x) = .12*10
.
.1x +1.5 - .15x = 1.2,,,,,combine like terms
.
-.05x = -.3
.
x=6 kg,,,,,,,,,,,,,,y=10-6=4 kg
.
.
check,,,,,.1(6) +.15(4) = .6 + .6 =1.2 = .12*10 ,,,,,,ok